# Existence and qualitative behavior of solutions of abstract differential-algebraic equations

### v1

@@ 1 Introduction
In the present paper, we study both the existence and uniqueness of solutions and their behavior with increasing time (boundedness and blow-up of solutions). From a practical point of view, it is important not only to prove that a solution exists on some local interval, but also to determine the maximal interval of its existence. If the solution is global (it exists on an arbitrarily large time interval), then the corresponding real-world system will operate over a sufficiently long duration without interruption. If the solution is ‘‘blow-up’’ in finite time (it exists on a finite time interval and is unbounded), then the operation of the corresponding real system will be interrupted. For example, the disruption of a semiconductor is described as the blow-up of a solution of an appropriate initial value or initial boundary value problem.
DAEs and ADAEs arise from the modelling of systems and processes in control problems, mechanics, robotics, gas industry, chemical kinetics, radio engineering, economics and other fields. It is well known that DAEs are used in modelling various objects whose structure is described by directed graphs, e.g., electrical, gas and neural networks. The application of DAEs is described in gas network theory, e.g., in [20, 10], and in electrical circuit theory is shown, e.g., in [33, 24, 6, 25, 9, 8, 22]. The DAEs of higher index arise, for example, in electrical engineering, gas industry, robotics, control problems and chemical kinetics (see, e.g., [2, 33, 24, 6, 25, 22]).
@@
where $A$ and $B$ are closed linear operators from $X$ into $Y$ with domains $D_{A}$ and $D_{B}$ respectively, $D=D_{A}\cap D_{B}\neq\{0\}$ is a lineal (linear manifold), $X$ and $Y$ are Banach spaces (BSs), $D_{A}$ and $D_{B}$ are dense in $X$ (i.e., $\overline{D_{A}}=X$, $\overline{D_{B}}=X$), $\mathop{\mathrm{range}}(A)=\overline{\mathop{\mathrm{range}}(A)}$ and $f\in C({\mathbb{R}}_{+}\times D,Y)$, ${\mathbb{R}}_{+}=[0,\infty)$. The operator $A$, as well as $B$, can be noninvertible (degenerate). Therefore, the ADAE (1.1) is also called an degenerate (or abstract) differential equation and degenerate (or abstract) evolution equation. According to the DAE terminology, the equation of the type (1.1) is called a semilinear ADAE.
@@
For linear ADAEs in Hilbert spaces, the condition for the existence and uniqueness of solutions on a finite interval in $\mathbb{R}$} have been presented, e.g., in [22, Section 12.3.3] (see also references on p. 577). The Galerkin method has been used to obtain these results. The existence and uniqueness of local solutions of certain semi-explicit and semilinear ADAEs of index 2 (in the sense that a stable semi-discretization leads to the finite-dimensional DAE of tractability index 2) has been studied in [14].
@@
  • $\overline{D}$ denotes the closure of a set $D$; $\partial D$ is the boundary of $D$; $D^{c}$ is the complement of $D$;
@@
  • $\mathcal{A}^{(-1)}$ is the semi-inverse operator of an operator $\mathcal{A}$  ($A^{-1}$ is the inverse operator of AA) (the definition of a semi-inverse operator can be found in [5]; see also [21] where such an operator is called an algebraic generalized inverse);
@@
  • if $V$ is a continuously differentiable functional, defined in a Banach space $X$, then the gradient of $V$ is denoted by $\partial_{x}V(x)$ or $\mathop{\mathrm{grad}}V(x)$ ($\mathop{\mathrm{grad}}V\colon X\to X^{*}$);
  • $\int\limits^{\infty}U(u)du=\infty$  ($\infty$}) means that the integral $\int\limits_{u_{0}}^{\infty}U(u)du=\infty$, i.e., does not converge ($\infty$}, i.e., converges) for any $u_{0}>0$;
@@ 2 Index of a regular pencil of operators, direct decompositions of spaces and the associated projectors
defined on $Y$ is called a regular point of the pencil $P(\lambda)$. The operator $R(\lambda)$ is called the resolvent of $P(\lambda)$. By $\varrho=\varrho(A,B):=\{\lambda\in{\mathbb{C}}\mid\exists\,(\lambda A+B)^{-1}\in\mathrm{L}(Y,X)\}$ (cf. [25]) we denote the set of the regular points $\lambda$ defined above. The pencil $P(\lambda)$ is said to be regular if $\varrho\neq\emptyset$ and to be singular (nonregular, irregular) if $\varrho=\emptyset$. The set $\varrho$ is open, and the resolvent as the operator function $R\colon\varrho\to\mathrm{L}(Y,X)$ is holomorphic on $\varrho$. Indeed, for each $\lambda_{*}\in\varrho$, it follows from the relation $R(\lambda)=R(\lambda_{*})[I_{Y}-(\lambda_{*}-\lambda)AR(\lambda_{*})]^{-1}$ and inclusion $AR(\lambda_{*})\in\mathrm{L}(Y)$ that $R(\lambda)=\sum\limits_{k=0}^{\infty}(\lambda_{*}-\lambda)^{k}R(\lambda_{*})\big(AR(\lambda_{*})\big)^{k}$ for $|\lambda-\lambda_{*}|<\big\|AR(\lambda_{*})\big\|_{\mathrm{L}(Y)}^{-1}$ (cf. [25]). Moreover, the point $\lambda=\infty$ is said to be a regular point of $P(\lambda)$ if in some neighborhood of the infinity ($|\lambda|>M$) all the points $\lambda$ are regular and $\|R(\lambda)\|\leq C|\lambda|^{-1}$ ($|\lambda|>M$) [25].
@@ 2.2 Direct decompositions of the operator domains and spaces and the associated projectors
Let $P(\lambda)$ be a regular pencil of index $\nu$, i.e., the above conditions (a) and (b) (or (2.2)) hold. Then by [25, Theorem 2.1] or [33, Section 2.3.1], there exists the pair of mutually complementary projectors $P_{k}\colon D\to D_{k}$ (by construction, $P_{k}$, $k=1,2$, have the domain $D_{A}$, $P_{1}D_{A}=P_{1}D$,  $P_{i}P_{j}x=\delta_{ij}P_{i}x$, $(P_{1}+P_{2})x=x$, $x\in D_{A}$) and the pair of mutually complementary projectors $Q_{k}\colon Y\to Y_{k}$ ($Q_{i}Q_{j}=\delta_{ij}Q_{i}$, $Q_{1}+Q_{2}=I_{Y}$), $k=1,2$, which generate the decompositions of $D$ and $Y$ into the direct sums
@@
are such that there exist $A_{1}^{-1}\in\mathrm{L}(Y_{1},\overline{D_{1}})$ and $B_{2}^{-1}\in\mathrm{L}(Y_{2},\overline{D_{2}})$. The projectors can be constructively determined by using contour integration [25, 33] or residues [8, 9]. Here $Q_{k}\in\mathrm{L}(Y)$, $k=1,2$. In addition, if we use the norm $\|\cdot\|_{X_{A}}$ of the space $X_{A}$ (see above), then $P_{k}\in\mathrm{L}(X_{A})$, $k=1,2$, and they generate the direct decomposition $X_{A}=D_{1}\mathop{\dot{+}}P_{2}D_{A}$, where $D_{1}=P_{1}D=P_{1}D_{A}$ (see [25]).
@@
In what follows, we suppose that $D_{B}\supseteq D_{A}$. Then $D=D_{A}$ and $\overline{D}=X$.
Denote $D_{20}=\ker A$. Assume that $\dim\ker A=n<\infty$ and denote by $\{\varphi_{1}^{1},...,\varphi_{n}^{1}\}$ a basis of $D_{20}$.
Note that if the operator $AR(\lambda_{*})$ is completely continuous (compact) for some $\lambda_{*}\in\varrho$, then the resolvent $R(\lambda)$ is completely continuous for each $\lambda\in\varrho$ (since $R(\lambda)=R(\lambda_{*})+(\lambda_{*}-\lambda)R(\lambda)AR(\lambda_{*})$), and the same holds for the resolvent $\widehat{R}(\mu)$  ($\mu=\lambda^{-1}$). Therefore, in this case the number of linearly independent eigenvectors of the operator $\widehat{P}(\mu)$ that correspond to the eigenvalue $\mu=0$ is finite [16] (recall that $\varphi$ is an eigenvector of $\widehat{P}(\mu)$, associated with the eigenvalue $\mu=0$, if $\widehat{P}(0)\varphi=A\varphi=0$ ), and hence $\dim\ker A<\infty$.
Using [16], we obtain that there exists a canonical system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ ($1\leq m_{i}\leq\nu$) of eigenvectors and adjoined vectors (or root vectors) of the pencil $\widehat{P}(\mu)=A+\mu B$ that correspond to the eigenvalue $\mu=0$, for which the vectors satisfy the equalities
@@
where $\nu$ is the index of $P(\lambda)$. The number $m_{i}-1$ is called the order of the adjoined vector $\varphi_{i}^{m_{i}}$, and $m_{i}$ is called the multiplicity of the eigenvector $\varphi_{i}^{1}$. If the eigenvector $\varphi_{i}^{1}$ does not have adjoined vectors (in this case, $\langle B\varphi_{i}^{1},q_{i}^{1}\rangle=1$, where $q_{i}^{1}\in\ker A^{*}$), then its multiplicity $m_{i}=1$ and the corresponding chain of root vectors consists only of $\varphi_{i}^{1}$. The vectors $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ and $\{B\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ form the bases of $D_{2}$ and $Y_{2}$, respectively (note that $D_{2}$ is closed), and the set $D_{2}$ is called the root subspace of $\widehat{P}(\mu)$ that correspond to $\mu=0$, and $\sum_{i=1}^{n}m_{i}$ is the root number of $\widehat{P}(\mu)$ (cf. [12]). Thus, the projectors $P_{k}$, $Q_{k}$, $k=1,2$, can be obtained by the formulas:
@@
where  $p_{i}^{j}=B^{*}q_{i}^{j}\in X^{*}$},  $j=1,...,m_{i}$}, i=1,,n{i=1,...,n}, $x\in X$, $y\in Y$, and the bounded linear functionals $q_{i}^{j}\in Y^{*}$ are chosen such that
@@
If $m_{i}=1$ (for some $i$), then (2.8) becomes $A^{*}q_{i}^{m_{i}}=0$. Note that since $\langle B\varphi_{i}^{m_{i}},q_{i}^{m_{i}}\rangle=1$, then $B\varphi_{i}^{m_{i}}\not\in\mathop{\mathrm{range}}(A)$, i=1,,ni=1,...,n. The formulas (2.7) and relations (2.9), (2.8) follow from the required properties of the operators $P_{k}$, $Q_{k}$, $k=1,2$, namely, they are the pairs of mutually complementary projectors and $\{D_{k},Y_{k}\}$, where $D_{k}=P_{k}D$ and $Y_{k}=Q_{k}Y$, are invariant under $A$, $B$, i.e.,
@@
The projectors (2.7) generate the decompositions of $X$, $D$ (recall that $D=D_{A}$ by assumption) and $Y$ into the direct sums
$\begin{split}&X=X_{1}\mathop{\dot{+}}X_{2},\quad X_{k}=P_{k}X,\quad D=D_{1}\mathop{\dot{+}}D_{2},\quad D_{k}=P_{k}D,\quad D_{2}=X_{2},\quad D_{1}=X_{1}\cap D,\\ &Y=Y_{1}\mathop{\dot{+}}Y_{2},\quad Y_{k}=Q_{k}Y\quad(Y_{2}=BD_{2}),\quad k=1,2,\end{split}$ (2.10)
@@
and $\{\hat{q}_{i}^{j}=q_{i}^{m_{i}+1-j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ is a canonical system of eigenvectors and adjoined vectors of the pencil $A^{*}+\mu B^{*}$ that correspond to the eigenvalue $\mu=0$} (this system is determined uniquely when specifying the system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$}) [16]. Therefore, $\{\hat{q}_{i}^{1}=q_{i}^{m_{i}}\}_{i=1,...,n}$} forms the basis of $\ker A^{*}$ and hence $\dim\ker A^{*}=n$}.
For the pencil $P(\lambda)$ of index 2, a similar construction of the projectors was used in [27] with the reference to [18]. According to [32], the system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ can be referred to as (B)(-B)-Jordan set of the operator $A$.
Recall that $D_{20}=\ker A=\mathop{\mathrm{span}}\{\varphi_{i}^{1}\}_{i=1,...,n}$}. Note that if $P(\lambda)$ has index 1, then $D_{2}=D_{20}=\ker A$.
@@
Notice that $D_{2\Sigma}=D_{2}\setminus D_{20}=\mathop{\dot{+}}\limits_{s=1}^{\nu-1}D_{2s}$},  $Y_{2*}=\mathop{\dot{+}}\limits_{s=0}^{\nu-1}Y_{2s}^{(s+1)}=\mathop{\dot{+}}\limits_{s=0}^{\nu-1}BD_{2s}^{(s+1)}$},   $Y_{2\Sigma}=\mathop{\mathrm{range}}(A_{2})$} and $Y_{1}=\mathop{\mathrm{range}}(A_{1})$}, where $A_{k}=A\big|_{D_{k}}\colon D_{k}\to Y_{k}$}, $k=1,2$ (they were introduced in (2.5)), and, accordingly, $\mathop{\mathrm{range}}(A)=Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$. Here it is supposed that the pencil index $\nu\geq 2$}. If $\nu=1$}, then $D_{2}=\ker A=\mathop{\mathrm{span}}\{\varphi_{i}^{1}\}_{i=1,...,n}$},  $Y_{2}=BD_{2}$} and $Y_{1}=\mathop{\mathrm{range}}(A)$}. As mentioned above, the operators $A_{1}=A\big|_{D_{1}}$ and $B_{2}=B\big|_{D_{2}}$ have the inverses
@@
where $x=x_{1}+x_{2\Sigma}+x_{20}$ (see (2.12)). The system (2.14), (2.15) is equivalent to the equation

### v2

@@ 1 Introduction
In the present paper, we study both the existence and uniqueness of solutions and their behavior with increasing time (boundedness and blow-up of solutions). From a practical point of view, it is important not only to prove that a solution exists on some local interval, but also to determine the maximal interval of its existence. If the solution is global (it exists on an arbitrarily large time interval), then the corresponding real-world system will operate over a sufficiently long duration without interruption. If the solution is “blow-up” in finite time (it exists on a finite time interval and is unbounded), then the operation of the corresponding real system will be interrupted. For example, the disruption of a semiconductor is described as the blow-up of a solution of an appropriate initial value or initial boundary value problem.
DAEs and ADAEs arise from the modelling of systems and processes in control problems, mechanics, robotics, gas industry, chemical kinetics, radio engineering, economics and other fields. It is well known that DAEs are used in modelling various objects whose structure is described by directed graphs, e.g., electrical, gas and neural networks. The application of DAEs is described in gas network theory, e.g., in [5, 3, 11], and in electrical circuit theory is shown, e.g., in [32, 23, 7, 24, 10, 9, 3, 21]. The DAEs of higher index arise, for example, in electrical engineering, gas industry, robotics, control problems and chemical kinetics (see, e.g., [3, 32, 23, 7, 24, 21]).
@@
where $A$ and $B$ are closed linear operators from $X$ into $Y$ with domains $D_{A}$ and $D_{B}$ respectively, $D=D_{A}\cap D_{B}\neq\{0\}$ is a lineal (linear manifold), $X$ and $Y$ are Banach spaces (BSs), $D_{A}$ and $D_{B}$ are dense in $X$ (i.e., $\overline{D_{A}}=\overline{D_{B}}=X$), $\mathop{\mathrm{range}}(A)=\overline{\mathop{\mathrm{range}}(A)}$ and $f\in C({\mathbb{R}}_{+}\times D,Y)$, ${\mathbb{R}}_{+}=[0,\infty)$. The operator $A$, as well as $B$, can be noninvertible (degenerate). Therefore, the ADAE (1.1) is also called a degenerate (or abstract) differential equation and degenerate (or abstract) evolution equation. According to the DAE terminology, the equation of the type (1.1) is called a semilinear ADAE.
@@
For linear ADAEs in Hilbert spaces, the condition for the existence and uniqueness of solutions on a finite interval in $\mathbb{R}$} have been presented, e.g., in [21, Section 12.3.3] (see also references on p. 577). The Galerkin method has been used to obtain these results. The existence and uniqueness of local solutions of certain semi-explicit and semilinear ADAEs of index 2 (in the sense that a stable semi-discretization leads to the finite-dimensional DAE of tractability index 2) has been studied in [14].
@@
  • $\overline{D}$ denotes the closure of a set $D$; $\partial D$ is the boundary of $D$; $D^{c}$ is the complement of $D$; $D_{A}$ is the domain of a linear operator $A$;
@@
  • $\mathcal{A}^{(-1)}$ is the semi-inverse operator of an operator $\mathcal{A}$  ($A^{-1}$ is the inverse operator of AA) (the definition of a semi-inverse operator can be found in [6]; see also [20] where such an operator is called an algebraic generalized inverse and, in Banach spaces, a topological generalized inverse);
@@
  • if $V$ is a continuously differentiable functional, defined in a Banach space $X$, then the gradient of $V$ is denoted by $\partial_{x}V(x)$ or $\mathop{\mathrm{grad}}V(x)$;
  • $\int\limits^{\infty}U(u)du=\infty$  ($\infty$}) means that the integral $\int\limits_{u_{0}}^{\infty}U(u)du=\infty$  ($\infty$}), i.e., does not converge (i.e., converges) for any $u_{0}>0$;
@@ 2 Index of a regular pencil of operators, direct decompositions of spaces and the associated projectors
defined on $Y$ is called a regular point of the pencil $P(\lambda)$. The operator $R(\lambda)$ is called the resolvent of $P(\lambda)$. By $\varrho=\varrho(A,B):=\{\lambda\in{\mathbb{C}}\mid\exists\,(\lambda A+B)^{-1}\in\mathrm{L}(Y,X)\}$ (cf. [24]) we denote the set of the regular points $\lambda$ defined above. The pencil $P(\lambda)$ is said to be regular if $\varrho\neq\emptyset$ and to be singular (nonregular, irregular) if $\varrho=\emptyset$. The set $\varrho$ is open, and the resolvent as the operator function $R\colon\varrho\to\mathrm{L}(Y,X)$ is holomorphic on $\varrho$. Indeed, for each $\lambda_{*}\in\varrho$, it follows from the relation $R(\lambda)=R(\lambda_{*})\big[I_{Y}-(\lambda_{*}-\lambda)AR(\lambda_{*})\big]^{-1}$ and inclusion $AR(\lambda_{*})\in\mathrm{L}(Y)$ that $R(\lambda)=\sum\limits_{k=0}^{\infty}(\lambda_{*}-\lambda)^{k}R(\lambda_{*})\big(AR(\lambda_{*})\big)^{k}$ for $|\lambda-\lambda_{*}|<\big\|AR(\lambda_{*})\big\|_{\mathrm{L}(Y)}^{-1}$ (cf. [24]). Moreover, the point $\lambda=\infty$ is said to be a regular point of $P(\lambda)$ if in some neighborhood of the infinity ($|\lambda|>M$) all the points $\lambda$ are regular and $\|R(\lambda)\|\leq C|\lambda|^{-1}$ for $|\lambda|>M$ [24].
@@ 2.2 Direct decompositions of the operator domains and spaces and the associated projectors
Let $P(\lambda)$ be a regular pencil of index $\nu$, i.e., the above conditions (a) and (b) (or (2.2)) hold. Then by [24, Theorem 2.1] or [32, Section 2.3.1], there exists the pair of mutually complementary projectors $P_{k}\colon D\to D_{k}$ (by construction, $P_{k}$, $k=1,2$, have the domain $D_{A}$, $P_{1}D_{A}=P_{1}D$,  $P_{i}P_{j}x=\delta_{ij}P_{i}x$, $(P_{1}+P_{2})x=x$, $x\in D_{A}$) and the pair of mutually complementary projectors $Q_{k}\colon Y\to Y_{k}$ ($Q_{i}Q_{j}=\delta_{ij}Q_{i}$, $Q_{1}+Q_{2}=I_{Y}$), $k=1,2$, which generate the decompositions of $D$ and $Y$ into the direct sums
@@
are such that there exist $A_{1}^{-1}\in\mathrm{L}(Y_{1},\overline{D_{1}})$ and $B_{2}^{-1}\in\mathrm{L}(Y_{2},\overline{D_{2}})$. The projectors can be constructively determined by using contour integration [24, 32] or residues [9, 10]. Here $Q_{k}\in\mathrm{L}(Y)$, $k=1,2$. In addition, if we use the norm $\|\cdot\|_{X_{A}}$ of the space $X_{A}$ defined above, then $P_{k}\in\mathrm{L}(X_{A})$, $k=1,2$, and they generate the direct decomposition $X_{A}=D_{1}\mathop{\dot{+}}P_{2}D_{A}$, where $D_{1}=P_{1}D=P_{1}D_{A}$ (see [24]).
@@
In what follows, we suppose that $\overline{D}=X$. Denote $D_{20}=\ker A\cap D$. In particular, if $D_{B}\supseteq D_{A}$, then $\overline{D}=X$ and $D_{20}=\ker A$.
Assume that $\dim D_{20}=n$ and denote by $\{\varphi_{1}^{1},...,\varphi_{n}^{1}\}$ a basis of $D_{20}$. Thus, $D_{20}=\mathop{\mathrm{span}}\{\varphi_{i}^{1}\}_{i=1,...,n}$} is closed.
Note that if the operator $AR(\lambda_{*})$ is completely continuous (compact) for some $\lambda_{*}\in\varrho$, then the resolvent $R(\lambda)$ is completely continuous for each $\lambda\in\varrho$ (since $R(\lambda)=R(\lambda_{*})+(\lambda_{*}-\lambda)R(\lambda)AR(\lambda_{*})$), and the same holds for the resolvent $\widehat{R}(\mu)$  ($\mu=\lambda^{-1}$). Therefore, in this case the number of linearly independent eigenvectors of the operator $\widehat{P}(\mu)$ that correspond to the eigenvalue $\mu=0$ is finite [16] (recall that $\varphi\in D$ is an eigenvector of $\widehat{P}(\mu)$, associated with the eigenvalue $\mu=0$, if $\widehat{P}(0)\varphi=A\varphi=0$ ), and hence $\dim D_{20}<\infty$.
Using [16], we obtain that there exists a canonical system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ ($1\leq m_{i}\leq\nu$) of eigenvectors and adjoined vectors (or root vectors) of the pencil $\widehat{P}(\mu)=A+\mu B$ that correspond to the eigenvalue $\mu=0$, for which the vectors $\varphi_{i}^{j}\in D$ satisfy the equalities
@@
where $\nu$ is the index of $P(\lambda)$. The number $m_{i}-1$ is called the order of the adjoined vector $\varphi_{i}^{m_{i}}$, and $m_{i}$ is called the multiplicity of the eigenvector $\varphi_{i}^{1}$. If the eigenvector $\varphi_{i}^{1}$ does not have adjoined vectors (in this case, $\langle B\varphi_{i}^{1},q_{i}^{1}\rangle=1$, where $q_{i}^{1}\in\ker A^{*}$), then its multiplicity $m_{i}=1$ and the corresponding chain of root vectors consists only of $\varphi_{i}^{1}$. The vectors $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ and $\{B\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ form the bases of $D_{2}$ and $Y_{2}$, respectively, and the set $D_{2}$ (note that $D_{2}$ is closed) is called the root subspace of $\widehat{P}(\mu)$ that correspond to $\mu=0$, and $\sum_{i=1}^{n}m_{i}$ is the root number of $\widehat{P}(\mu)$ (cf. [13]). Thus, the projectors $P_{k}$, $Q_{k}$, $k=1,2$, can be obtained by the formulas:
@@
where  $p_{i}^{j}=B^{*}q_{i}^{j}\in X^{*}$},  $j=1,...,m_{i}$}, i=1,,n{i=1,...,n}, $x\in X$, $y\in Y$, and the bounded linear functionals $q_{i}^{j}\in D^{*}=D_{A^{*}}\cap D_{B^{*}}$ are chosen such that
@@
If $m_{i}=1$ (for some $i$), then (2.8) becomes $A^{*}q_{i}^{m_{i}}=0$. Note that since $\langle B\varphi_{i}^{m_{i}},q_{i}^{m_{i}}\rangle=1$  ($q_{i}^{m_{i}}\in\ker A^{*}$), then $B\varphi_{i}^{m_{i}}\not\in\mathop{\mathrm{range}}(A)$, i=1,,ni=1,...,n. The formulas (2.7) and relations (2.9), (2.8) follow from the required properties of the operators $P_{k}$, $Q_{k}$, $k=1,2$, namely, they are the pairs of mutually complementary projectors and $\{D_{k},Y_{k}\}$, where $D_{k}=P_{k}D$ and $Y_{k}=Q_{k}Y$, are invariant under $A$, $B$, i.e.,
@@
and $\{\hat{q}_{i}^{j}=q_{i}^{m_{i}+1-j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ is a canonical system of eigenvectors and adjoined vectors of the pencil $A^{*}+\mu B^{*}$ that correspond to the eigenvalue $\mu=0$} (this system is determined uniquely when specifying the system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$}) [16]. Therefore, $\{\hat{q}_{i}^{1}=q_{i}^{m_{i}}\}_{i=1,...,n}$} forms the basis of $D_{20}^{*}=\ker A^{*}\cap D^{*}$ and hence $\dim D_{20}^{*}=n$}. If $D_{B}\supseteq D_{A}$, then $\{\hat{q}_{i}^{1}\}_{i=1,...,n}$} is the basis of $\ker A^{*}$ and $\dim\ker A^{*}=n$}.
The projectors (2.7) generate the decompositions of $D$, $X$ and $Y$ into the direct sums
$\begin{split}&D=D_{1}\mathop{\dot{+}}D_{2},\quad D_{k}=P_{k}D,\quad X=X_{1}\mathop{\dot{+}}X_{2},\quad X_{k}=P_{k}X,\quad D_{2}=X_{2},\quad D_{1}=X_{1}\cap D,\\ &Y=Y_{1}\mathop{\dot{+}}Y_{2},\quad Y_{k}=Q_{k}Y\quad(Y_{2}=BD_{2}),\quad k=1,2,\end{split}$ (2.10)
For the pencil $P(\lambda)$ of index 2, a similar construction of the projectors was used in [26] with the reference to [18]. If $D_{B}\supseteq D_{A}$, then, according to [31], the system $\{\varphi_{i}^{j}\}_{i=1,...,n}^{j=1,...,m_{i}}$ can be referred to as (B)(-B)-Jordan set of the operator $A$.
Recall that $D_{20}=\ker A\cap D=\mathop{\mathrm{span}}\{\varphi_{i}^{1}\}_{i=1,...,n}$}. Note that if $P(\lambda)$ has index 1, then $D_{2}=D_{20}$ and $Y_{1}=\mathop{\mathrm{range}}\big(A\big|_{D}\big)$ (if in addition $D_{B}\supseteq D_{A}$, then $D_{2}=\ker A$ and $Y_{1}=\mathop{\mathrm{range}}(A)$ ).
@@
Notice that $D_{2\Sigma}=D_{2}\setminus D_{20}=\mathop{\dot{+}}\limits_{s=1}^{\nu-1}D_{2s}$},  $Y_{2*}=\mathop{\dot{+}}\limits_{s=0}^{\nu-1}Y_{2s}^{(s+1)}=\mathop{\dot{+}}\limits_{s=0}^{\nu-1}BD_{2s}^{(s+1)}$},   $Y_{2\Sigma}=\mathop{\mathrm{range}}(A_{2})$} and $Y_{1}=\mathop{\mathrm{range}}(A_{1})$}, where $A_{k}=A\big|_{D_{k}}\colon D_{k}\to Y_{k}$}, $k=1,2$ (they were introduced in (2.5)), and, accordingly, $\mathop{\mathrm{range}}\big(A\big|_{D}\big)=Y_{1}\mathop{\dot{+}}Y_{2\Sigma}$. Here it is supposed that the pencil index $\nu\geq 2$}. If $\nu=1$}, then $D_{2}=D_{20}=\mathop{\mathrm{span}}\{\varphi_{i}^{1}\}_{i=1,...,n}$},  $Y_{2}=BD_{2}$} and $Y_{1}=AD$}. As mentioned above, the operators $A_{1}=A\big|_{D_{1}}$ and $B_{2}=B\big|_{D_{2}}$ have the inverses
@@
where $x=x_{1}+x_{2\Sigma}+x_{20}\in D$ (see (2.12)). The system (2.14), (2.15) is equivalent to the equation